The force between a 100 kg man and the Earth is 980 N. How close must two protons (1.6 X 10-19 C) be to generate the same force?


1
Expert's answer
2022-04-14T15:45:43-0400

"d=\sqrt{\frac {kp^2}F}=0.5\cdot 10^{-15}~m."